OpenStudy (anonymous):

FIND THE DERIVATIVE OF SQUARE ROOT OF 3+4X

OpenStudy (tkhunny):

Have you considered the Chain Rule?

OpenStudy (anonymous):

d/dx √(3+4x) = f'(g(x))*g'(x) [chain rule] 1/2 (3+4x)^(1/2 -1) * 4 g'(x)=g'(3+4x) = 0+4=4 (1/2)(3+4x)^(-1/2) *4 2(3+4x)^(-1/2) 0r 2/√(2+4x)

OpenStudy (anonymous):

2 / radical ( 3+4x)